# Strategic Tradeoffs Between Humans and AI in Multi-Agent Bargaining

### v2

@@ Abstract
Coordination tasks traditionally performed by humans are increasingly being delegated to autonomous agents. As this pattern progresses, it becomes critical to evaluate not only these agents’ performance but also the processes through which they negotiate in dynamic, multi-agent environments. Furthermore, different agents exhibit distinct advantages: traditional statistical agents, such as Bayesian models, may excel under well-specified conditions, whereas large language models (LLMs) can generalize across contexts. In this work, we compare humans (N = 216), LLMs (GPT-4o, Gemini 1.5 Pro), and Bayesian agents in a dynamic negotiation setting that enables direct, identical-condition comparisons across populations, capturing both outcomes and behavioral dynamics. Bayesian agents extract the highest surplus through aggressive optimization, at the cost of frequent trade rejections. Humans and LLMs can achieve similar overall surplus, but through distinct behaviors: LLMs favor conservative, concessionary trades with few rejections, while humans employ more strategic, risk-taking, and fairness-oriented behaviors. Thus, we find that performance parity—a common benchmark in agent evaluation—can conceal fundamental differences in process and alignment, which are critical for practical deployment in real-world coordination tasks.
@@ 1 Introduction
Historically, bespoke statistical models have excelled in structured economic exchanges with well-defined action spaces (Sun and Müller, 2013; Dehghanpour et al., 2016). More recently, the generalizability and flexibility of large language models (LLMs) have enabled delegation in unstructured contexts: negotiating supply chain contracts (Van Hoek et al., 2022), coordinating international trading (Corvin, 2024), and facilitating online commerce transactions (PYMNTS, 2025; Gaarlandt et al., 2025). Unlike more structured models, LLMs can generalize under uncertainty and partial information, reducing the need for task-specific customization. As AI agents grow more sophisticated and are entrusted with higher-stakes decisions, understanding how different types of agents behave in dynamic, multi-agent settings becomes essential. In particular, navigating the trade-offs between agent types—not only in terms of negotiated outcomes, but also in their alignment with human strategies and norms—is critical for responsible delegation.
@@
We address this challenge by introducing a novel multi-player bargaining game that enables direct comparison of populations (e.g., humans, LLMs, and statistical models) under identical conditions. In our game, players receive endowments of colored chips with private valuations and negotiate trades to maximize their individual surplus. Bargaining is an ideal testbed for dynamic, multi-agent evaluations, as it inherently exposes tensions between social norms, individual preferences, and strategic reasoning, reflecting the complexities of many real-world environments. Moreover, from a theoretical perspective, bargaining under partial information admits no Pareto efficient, incentive-compatible solutions (Myerson and Satterthwaite, 1983), forcing agents to navigate trade-offs between individual gain and mutually beneficial cooperation.
We conduct an empirical study of our game across three populations: (i) human participants (N = 216), (ii) LLM-based agents, and (iii) specialized Bayesian agents, all facing identical randomized endowments. The Bayesian agents achieved the highest surplus, proposing aggressive bargains at the cost of frequent rejections. LLMs matched human surplus in certain game conditions but exhibited fundamentally different behaviors: where humans made fairness-driven concessions, LLMs favored conservative, low-variance trades with high acceptance rates.
More broadly, these findings expose a limitation of outcome-based evaluation: aggregate surplus can mask important differences in behavior and alignment. As AI systems increasingly mediate real-world negotiations, it becomes essential to assess not just what agents achieve, but how they navigate trade-offs, adjust to strategic dynamics, and engage with social norms.
@@ 2 Related work
AI agents have excelled in well-defined tasks with explicit objectives, preferences, and constrained action spaces (Sun and Müller, 2013; Dehghanpour et al., 2016). Recent advances in LLMs have broadened the scope of delegation to include tasks requiring broader social reasoning, such as preference elicitation, strategic pricing, natural-language negotiation, and alliance formation in multi-agent games like Diplomacy (Fish et al., 2024; Soumalias et al., 2025; Tessler et al., 2024; Vaccaro et al., 2025; Lubars and Tan, 2019; Bakhtin et al., 2022; Kramár et al., 2022; Deng et al., 2024). Beyond social reasoning, LLMs exhibit emergent capabilities in economic computation and human alignment, replicating behaviors in bargaining, auctions, and other economic games (Horton, 2023; Aher et al., 2022; Argyle et al., 2022), and reflecting human subgroup-specific patterns when conditioned on demographic traits (Manning et al., 2024; Xie et al., 2024; Zhu et al., 2024; Park et al., 2023).
However, successful deployment depends not just on capabilities (Jahani et al., 2025), but also on user expectations, skills, and alignment with social norms; users may suboptimally delegate when system behavior violates these expectations (Mozannar et al., 2023; Agarwal et al., 2023; Mullainathan and Obermeyer, 2021; Palminteri et al., 2024; Kapania et al., 2022; Qian and Wexler, 2024). Addressing alignment gaps requires evaluating both capabilities and behavior in dynamic, social contexts—a need recognized by recent calls for rigorous methods to steer and assess agent behavior in strategic multi-agent settings (Raman et al., 2024; Goktas et al., 2025). To concretely examine these challenges, we study how humans, LLM-based agents, and Bayesian models behave in a dynamic bargaining task, offering a comparative lens on alignment in strategic social interactions.
@@ 3 The bargaining game
We introduce a multi-player bargaining game which captures strategic trade-offs, social signaling, and coordination challenges found in real-world negotiations. Figure 1 presents an overview of the game, in which three players aim to maximize individual surplus over nine turns. Each begins with ten chips of each color. Green chips act as a shared numeraire valued at \$0.50 by all; other chip colors (red, blue, purple) have private values, drawn uniformly between \$0.10 and \$1.00. Game complexity is varied by incrementally adding chip types: red, then blue, then purple, with green always present. [^1] We show in Appendix A.2 that decision complexity grows with the number of chip colors.
Figure 1: This provides a broad overview of the bargaining game. Left: Game setup. Center: Simplified version of gameplay. Right: The agent populations evaluated in this study.
@@ 4.2 Efficient allocations
Intuitively, in a centralized setting with full information, $\mathbf{A^{*}}$ represents a Pareto-efficient allocation that maximizes aggregate utility without making any agent worse off. While this optimal allocation $\mathbf{A^{*}}$ may not be unique, the maximum total welfare $\mathbf{w^{*}}$ achieved at the optimum is unique. This optimization problem can be solved efficiently using interior point methods. Its computational complexity is bounded by $O\bigl{(}(|I||G|)^{3.5}\bigr{)}$.
Note that this upper bound is unlikely to be achieved in practice, as no mechanism for decentralized bargaining with private information can guarantee both Pareto efficiency and incentive compatibility Myerson and Satterthwaite (1983). [^2] A proof that players of the bargaining game lack a dominant strategy is provided in AppendixA.1.
@@ 5 Experimental setup
In our design, human participants first completed games in groups of three. We then replicated each game (identical initial chip endowments) with LLM- and Bayesian- agent groups, enabling direct comparison across populations. [^3] For reproducibility, details on the human game interface are provided in Appendix B.1. Full LLM prompts are provided in Appendix G, and the Bayesian agent algorithms are detailed in Appendix A.3.
@@ 5.2 LLM agents
We created two types of agents. As our primary goal was to evaluate baseline performance under minimal optimization, we first constructed an out-of-box agent used simple prompts adapted from the human instructions. To assess the benefit of lightweight refinement, we also implemented a refined agent which generated candidate proposals and selected one using “type-2” reasoning, mimicking slower, more deliberate decision-making (Furniturewala et al., 2024).
To generalize our findings, we constructed LLM populations using two model families: OpenAI’s GPT-4o (OpenAI et al., 2024) and Google’s Gemini 1.5 Pro (Team et al., 2024), which were the latest full-capacity models from each family at time of writing. [^6] For robustness, we also benchmarked the latest lightweight models (Gemini 2.5 Flash, GPT-o4-mini), but they failed to execute basic surplus-maximizing trades. We therefore restrict our main analysis to full-capacity models; see Appendix F for additional context. We set prompt temperature to 0.5, following mid-range recommendations for structured reasoning tasks that require coherence and variability (Arora et al., 2024). Prompts were scaffolded using a chain-of-thought strategy through the open-source EDSL library (Wei et al., 2022; Horton et al., 2024).
@@ 5.3 Bayesian agents
The Bayesian agents were programmed specifically to play this bargaining game. Their strategy is formalized as follows. Each agent i forms a joint belief about the valuations of all other players, $\mathbf{v}_{-i}$. When agent i proposes a trade $(x_{g},y_{r})$ – offering $x$ chips of type $g$ in exchange for $y$ chips of type $r$ – they choose a trade to maximize their expected payoff:
@@ 6.1 Performance
Figure 2 shows surplus trajectories by agent types and game variation. In the upper right panel, we see that Bayesian agents reached 74% of the Pareto-efficient optimum in the two-chip game. Across all variations, Bayesian agents achieved the highest surplus. GPT-4o performed comparably to human participants; however, humans statistically outperformed Gemini 1.5 Pro. [^7] This may reflect the recency of such capabilities; Gemini 1.5 Pro was released three months prior to GPT-4o and represented an earlier generation of model development. Refined LLMs did not significantly impact surplus compared to out-of-the-box LLMs (statistical tests in Appendix E).
@@ 6.2.1 Trading patterns and volume dynamics
Figure 3: Trading patterns in the 3-chip game, visualizing (i) net surplus change for the proposer and (ii) trade ratio. Accepted trades in green, rejections in red. Marginal distributions across each axis are adjacent to the plots. The vertical dashed line marks zero net surplus (no net value created), and the horizontal solid line marks balanced exchange (1:1 ratio of chips). Figure values are in Table 5.
@@
Figure 3 plots individual trade proposals for the three-chip game for each agent type. Results for the two- and four-chip games are qualitatively similar (see Appendix F.2). The vertical balanced value line indicates proposals that generated no net surplus, while the horizontal balanced ratio line corresponds to 1:1 exchanges of chips. In the left panel, we see that the human proposals cluster around the balanced ratio line: that is, humans often offer a similar number of chips as they are requesting. Most trade proposals lie to the right of the balanced value line (i.e., provide positive surplus to the proposer), though some reduce total surplus.
@@
Both LLM populations propose trades near the vertical balanced net value line (yielding little total surplus), with no anchoring to balanced trade ratios. A noticeable vertical tail of proposals extends above the 1:1 balanced ratio line, indicating a tendency to offer more chips than requested (even up to 5:1). Many proposals also incur a net surplus loss.
@@
By construction, Bayesian agents strictly maximize surplus and avoid net-loss trades (no points left of the balanced value line). Their proposals reach deep into the positive surplus region. Rather than the 1:1 balanced ratio exhibited in human proposals, the Bayesian offers cluster below the parity line (ratio < 1), meaning they ask for more chips than they give. Proposal rejections are high.
@@ F.2 Supplementary trade space visualizations
Figure 13: Comparison of the trade space (4-Chip) between Human participants, LLM agents (GPT-4o and Gemini-1.5 Pro), and Bayesian agents.
@@ G.4 Refined Proposing Prompt
The refined prompt first generates multiple trades proposals and uses another LLM query to choose the best one among them.

### v3

@@ Abstract
As large language models (LLMs) are increasingly embedded in collaborative human activities such as business negotiations and group coordination, it becomes critical to evaluate both the performance gains they can achieve and how they interact in dynamic, multi-agent environments. Unlike traditional statistical agents such as Bayesian models, which may excel under well-specified conditions, large language models (LLMs) can generalize across diverse, real-world scenarios, raising new questions about how their strategies and behaviors compare to those of humans and other agent types. In this work, we compare outcomes and behavioral dynamics across humans (N = 216), LLMs (GPT-4o, Gemini 1.5 Pro), and Bayesian agents in a dynamic negotiation setting under identical conditions. Bayesian agents extract the highest surplus through aggressive optimization, at the cost of frequent trade rejections. Humans and LLMs achieve similar overall surplus, but through distinct behaviors: LLMs favor conservative, concessionary trades with few rejections, while humans employ more strategic, risk-taking, and fairness-oriented behaviors. Thus, we find that performance parity—a common benchmark in agent evaluation—can conceal fundamental differences in process and alignment, which are critical for practical deployment in real-world coordination tasks. By establishing foundational behavioral baselines under matched conditions, this work provides a baseline for future studies in more applied, variable-rich environments.
@@ 1 Introduction
Historically, bespoke statistical models have excelled in structured economic exchanges with well-defined action spaces (Sun and Müller, 2013; Dehghanpour et al., 2016). More recently, the generalizability and flexibility of large language models (LLMs) have enabled delegation in unstructured contexts: negotiating supply chain contracts (Van Hoek et al., 2022), coordinating international trading (Corvin, 2024), and facilitating online commerce transactions (PYMNTS, 2025; Gaarlandt et al., 2025). Unlike more structured models, LLMs can generalize under uncertainty and partial information, reducing the need for task-specific customization. As AI agents grow more sophisticated and are entrusted with higher-stakes decisions, understanding how different types of agents behave in dynamic, multi-agent settings is essential. In particular, navigating the trade-offs between agent types—not only in terms of negotiated outcomes, but also in their alignment with human strategies and norms—is critical for responsible delegation.
@@
We address this challenge by introducing a novel multi-player bargaining game that enables direct comparison of different agent types (e.g., humans, LLMs, and statistical models) under identical conditions. In our game, players receive endowments of colored chips with private valuations and negotiate trades to maximize their individual surplus. Bargaining is an ideal testbed for dynamic, multi-agent evaluations, as it inherently exposes tensions between social norms, individual preferences, and strategic reasoning, reflecting the complexities of real-world environments. Moreover, from a theoretical perspective, bargaining under partial information admits no Pareto efficient, incentive-compatible solutions (Myerson and Satterthwaite, 1983a), forcing agents to navigate trade-offs between individual gain and mutually beneficial cooperation.
We conduct an empirical study of our game across three types of agents: (i) human participants (N = 216), (ii) LLM-based agents, and (iii) specialized Bayesian agents, all facing identical randomized endowments. The Bayesian agents achieved the highest surplus, proposing aggressive bargains at the cost of frequent rejections. LLMs matched human surplus in certain game conditions but exhibited fundamentally different behaviors: where humans made fairness-driven concessions, LLMs favored conservative, low-variance trades with high acceptance rates.
More broadly, these findings expose a limitation of outcome-based evaluation: aggregate surplus can mask important differences in behavior and alignment. As AI systems increasingly mediate real-world negotiations, it becomes essential to assess not just what agents achieve, but how they navigate trade-offs, adjust to strategic dynamics, and engage with social norms. The primary contributions of this work are threefold:
  • A dynamic multi-agent environment for comparative evaluation. We design and implement a bargaining game that enables direct, identical-condition comparisons of negotiation strategies across different agent types, with game variations of increasing complexity. We accompany the platform with a linear programming formulation of the Pareto-optimal surplus, providing a theoretical benchmark for performance.
  • An empirical characterization of human, LLM, and Bayesian negotiation strategies. Through a study with 216 human participants (N=216), we reveal distinct behavioral profiles: Bayesian agents adopt aggressive, surplus-maximizing strategies with high rejection rates; humans exhibit strategic concessions and fairness norms; LLMs act as conservative agents that yield high rates of offer acceptance.
  • A framework to demonstrate how performance parity can conceal strategic misalignment. Our experimental benchmarks show that aggregate outcomes (i.e., total surplus) are insufficient for evaluation, as they mask fundamental differences in agent behavior. This finding highlights the need to move beyond performance-based metrics to assess process and alignment for the responsible deployment of agents in real-world coordination tasks.
@@ 2 Related work
Statistical agents have excelled in well-defined tasks with explicit objectives, preferences, and constrained action spaces (Sun and Müller, 2013; Dehghanpour et al., 2016). Recent advances in LLMs have broadened the scope of delegation to include tasks requiring broader social reasoning, such as preference elicitation, strategic pricing, natural-language negotiation, and alliance formation in multi-agent games like Diplomacy (Fish et al., 2024; Soumalias et al., 2025; Tessler et al., 2024; Vaccaro et al., 2025; Lubars and Tan, 2019; Bakhtin et al., 2022; Kramár et al., 2022; Deng et al., 2024). LLMs have exhibited emergent capabilities in economic computation and human alignment, replicating behaviors in bargaining, auctions, and other economic games (Horton, 2023; Aher et al., 2022; Argyle et al., 2022), and reflecting human subgroup-specific patterns when conditioned on demographic traits (Manning et al., 2024; Xie et al., 2024; Zhu et al., 2024; Park et al., 2023; Qian et al., 2025).
However, appropriate deployment and delegation depend not only on system capabilities (Jahani et al., 2025), but also on user expectations, skills, and alignment with social norms; users may suboptimally delegate when system behavior violates expectations (Mozannar et al., 2023; Agarwal et al., 2023; Mullainathan and Obermeyer, 2021; Palminteri et al., 2024; Kapania et al., 2022; Qian and Wexler, 2024). Addressing alignment gaps requires evaluating both capabilities and behaviors in dynamic, social contexts—a need recognized by recent calls for rigorous methods to steer and assess agent behavior in strategic multi-agent settings (Raman et al., 2024; Goktas et al., 2025). To concretely examine these challenges, we study how humans, LLM-based agents, and Bayesian models behave in a dynamic bargaining task, offering a comparative lens on alignment in strategic social interactions.
@@ 3 The bargaining game
3.1 Motivation from real-world negotiations
Negotiation remains one of humanity’s most fundamental coordination tasks Nash et al. (1950); Rubinstein (1982). Bilateral and multi-party bargaining appear in contexts such as three-team NBA tradesPelechrinis et al. (2023), B2B inventory swaps, and corporate barter networks Binmore (2007). These environments require participants to negotiate over bundles of heterogeneous assets with private valuations, making offers such as “X of A for Y of B (plus cash),” under real-world feasibility and time constraints.
Such bargaining is both strategically and communicatively demanding Farrell and Rabin (1996); Camerer (2003), motivating efforts to automate or augment the process. With advances in large language models (LLMs), it is now plausible to automate not just basic proposals, but also reasoning, persuasion, and adaptation in natural language. However, the strategic differences across LLM and human negotiation behavior remain underexplored; LLMs may diverge from human norms in ways that shape outcomes, affecting agreement rates, surplus allocation, and perceptions of fairness.
Studying these trade-offs in applied settings invites confounds; for example, if we were to evaluate the exchange of real commodities like secondhand goods in an online marketplace, participants’ heterogeneous preferences would confound direct comparison across agent types. To address this, we construct an abstract bargaining game that captures the essential structure of bargaining markets: fixed assets, a shared numeraire, private valuations, and an offer–response protocol Roth and Kagel (1995). The tensions embedded in our game—private information, strategic risk, and the trade-off between individual gain and mutual benefit—are fundamental to a wide range of real-world coordination tasks, from collective bargaining for labor contracts to negotiating international climate treaties Schelling (1980); Myerson and Satterthwaite (1983b). Therefore, our findings on agent behavior provide a foundational and generalizable diagnostic for understanding how these agents will likely act in more complex, unstructured interactions.
3.2 A stylized game
We introduce a multi-player bargaining game which captures strategic trade-offs, social signaling, and coordination challenges, influenced by real-world negotiations. Figure 1 presents an overview of the game, in which three players aim to maximize individual surplus over nine turns. Each begins with ten chips of each color. Green chips act as a shared numeraire valued at \$0.50 by all; other chip colors (red, blue, purple) have private values, drawn uniformly between \$0.10 and \$1.00. Game complexity is varied by incrementally adding chip types: red, then blue, then purple, with green always present. [^1] We show in Appendix A.2 that decision complexity grows with the number of chip colors.
Figure 1: This provides a broad overview of the bargaining game. Left: Game setup with initial chip counts and valuations. Center: Simplified version of gameplay (offer proposals and acceptances). Right: The types of agents empirically evaluated and simulated in this study.
@@ 4.2 Efficient allocations
Intuitively, in a centralized setting with full information, $\mathbf{A^{*}}$ represents a Pareto-efficient allocation that maximizes aggregate utility without making any agent worse off. While this optimal allocation $\mathbf{A^{*}}$ may not be unique, the maximum total welfare $\mathbf{w^{*}}$ achieved at the optimum is unique. This optimization problem can be solved efficiently using interior point methods. Its computational complexity is bounded by $O\bigl((|I||G|)^{3.5}\bigr)$. Note that this upper bound is unlikely to be achieved in practice, as no mechanism for decentralized bargaining with private information can guarantee both Pareto efficiency and incentive compatibility Myerson and Satterthwaite (1983a). [^2] A proof that players of the bargaining game lack a dominant strategy is provided in AppendixA.1.
@@ 5 Experimental setup
In our design, human participants first completed games in groups of three. Each group was randomly assigned to either the 2-chip, 3-chip, or 4-chip game variation, allowing us to evaluate performance and strategic trade-offs under increasing task complexity. We then replicated each game (identical initial chip endowments) with LLM- and Bayesian- agent groups, enabling direct comparison across the groups. [^3] For reproducibility, details on the human game interface are provided in Appendix B.1. Full LLM prompts are provided in Appendix G, and the Bayesian agent algorithms are detailed in Appendix A.3.
@@ 5.2 LLM agents
We developed two types of agents to evaluate the tradeoffs between simple prompting and chain-of-thought refinement. As our primary goal was to evaluate baseline performance under minimal optimization, we first constructed an out-of-box agent used simple prompts adapted from the human instructions. To assess the benefit of lightweight refinement, we also implemented a refined agent which generated candidate proposals and selected one using “type-2” reasoning, mimicking slower, more deliberate decision-making (Furniturewala et al., 2024). Each agent type was implemented across two model families for generalizability: OpenAI’s GPT-4o (OpenAI et al., 2024) and Google’s Gemini 1.5 Pro (Team et al., 2024), which were the latest full-capacity models from each family at time of writing. [^6] For robustness, we also benchmarked the latest lightweight models (Gemini 2.5 Flash, GPT-o4-mini), but they failed to execute basic surplus-maximizing trades. We therefore restrict our main analysis to full-capacity models; see Appendix F for additional context. Prompt temperature was set to 0.5, following mid-range recommendations for structured reasoning tasks that require coherence and variability (Arora et al., 2024). Prompts were scaffolded using a chain-of-thought strategy through the open-source EDSL library (Wei et al., 2022; Horton et al., 2024).
@@ 5.3 Bayesian agents
We include Bayesian players as a statistical benchmark: these agents represent the traditional, hard-coded approach to negotiation, explicitly programmed to optimize within the structure of this bargaining game. Each agent i forms a joint belief about the valuations of all other players, $\mathbf{v}_{-i}$. When agent i proposes a trade $(x_{g},y_{r})$ – offering $x$ chips of type $g$ in exchange for $y$ chips of type $r$ – they choose a trade to maximize their expected payoff:
@@ 6.1 Performance
Figure 2 shows surplus trajectories by agent type and game variation. Bayesian agents reached 74% of the Pareto-efficient optimum in the two-chip game (upper right panel). Across all variations, Bayesian agents achieved the highest surplus. GPT-4o performed comparably to human participants; however, humans statistically outperformed Gemini 1.5 Pro. [^7] This may reflect the recency of such capabilities; Gemini 1.5 Pro was released three months prior to GPT-4o and represented an earlier generation of model development. Refined LLMs did not significantly impact surplus compared to out-of-the-box LLMs (Appendix E).
@@ 6.2.1 Trading patterns and volume dynamics
Figure 3: Trading patterns in the 3-chip game, visualizing (i) net surplus change for the proposer and (ii) trade ratio. Accepted trades are shown in green, rejected trades in red. Marginal distributions across each axis are adjacent to the plots. The vertical dashed line marks zero net surplus (no net value created), and the horizontal solid line marks balanced exchange (1:1 ratio of chips). Figure values are provided in Table 5.
@@
Figure 3 plots individual trade proposals for the three-chip game for each agent type. Results for the two- and four-chip games are qualitatively similar (Appendix F.2). The vertical balanced value line indicates proposals that generated no net surplus, while the horizontal balanced ratio line corresponds to 1:1 exchanges of chips. In the left panel, we see that the human proposals cluster around the balanced ratio line: that is, humans often offer a similar number of chips as they are requesting. Most trade proposals lie to the right of the balanced value line (i.e., provide positive surplus to the proposer), though some reduce total surplus.
@@
Both LLM families propose trades near the vertical balanced net value line, yielding little total surplus with no anchoring to balanced trade ratios. A noticeable vertical tail of proposals extends above the 1:1 balanced ratio line, indicating a tendency to offer more chips than requested (even up to 5:1). Many proposals also incur a net surplus loss.
@@
By construction, Bayesian agents strictly maximize surplus (evidenced by positive surplus proposals) and avoid net-loss trades (no points left of the balanced value line). Rather than the 1:1 balanced ratio exhibited in human proposals, the Bayesian offers cluster below the parity line (ratio < 1), meaning they ask for more chips than they give. Proposal rejections are high.
Trade pattern comparisons.
@@ F.2 Supplementary trade space visualizations
Figure 13: Trade space comparison (4-chip) between humans, LLM agents (GPT-4o and Gemini-1.5 Pro), and Bayesian agents.
@@ G.4 Refined Proposing Prompt
The refined prompt first generates multiple trades proposals and uses another LLM query to choose among them.